# Ad-hoc Concept Forming in the Game Codenames as a Means for Evaluating Large Language Models

### v1

@@ 1 Introduction
We have implemented the game Codenames as a challenging means (already investigated in psycholinguistic studies, see below) for evaluating certain language-use capabilities. In this game, a first player needs to provide to a second player a clue which singles out certain words within a larger set of words given to both players (see Figure 6 below for examples).
@@ 4.2 In-depth Analysis
4.2.1 Number of Targets, Guesses & Revealed
In Figure 2, we present the average number of words targeted and words guesses by selected models. We can see that high-performing models such as o3-mini and Deepseek-r1 generate at least 1-2 more words as targets and guesses in the beginning. Targeting and guessing more words in a single turn is the standard strategy in Codenames to win Spendlove and Ventura (2022), especially needed when playing against the mock opponent, which reveals one word at each turn. Models tend to guess fewer words than were targeted. For instance, o3-mini on average targets more than four words but guessed considerably fewer for the first turn, unlike Deepseek-r1, which targets and guesses an almost equal number of words. In Figure 7, we included the average number of target, guessed and revealed (where the guess is team word) words per model. We can see that only Deepseek-r1 exceeds the threshold of more than two words (2.2), while the rest have close values (1.5-1.9). It indicates that all models guess wrong words by revealing words from the opponent team or distractors, or even assassin words.
4.2.2 Success, Lose & Aborted Lose Rates
@@ 4.2.3 Efficiency & Sensitivity
4.2.4 Typical Errors
To understand where models fail and how higher-performing models differ from lower ones, we analysed the most common errors, then categorised them and counted each occurrence, see Table 3.
@@ 4.2.4 Typical Errors
The differentiating factor in high-performing models is that hallucination and instruction following issues appear more rarely. For instance, the first error type, Target Hallucinated, refers to cases where Spymaster generates a clue and targets some words, but some of these do not exist on the board (as should be known to the model). In such cases, the GameMaster aborts that episode. Similarly, Guess Hallucinated is an error that occurs on the Field Operative side where it guesses a word that does not exist on the board. Mostly, Llama-3.1-405B and Llama-3.1-8B have another issue with guessing the correct number of words that the Spymaster indicates. They tend to guess more than the number of target words (note here: models can guess less but not more than target words). Lastly, the common issue, Guess is Clue, with low-performing models is that the guessed word is the same as the clue in many cases. It shows a lack of pragmatic reasoning for choosing unrevealed candidate words from available ones on a board. In all of these cases and some minor ones, e.g. tags such as “CLUE:”, “TARGETS:”, “GUESS:” are omitted, the GameMaster aborts the game because the rules are not followed. Such instruction-following issues happen mostly with Deepseek-r1. See Table 5 for the complete list.

### v2

@@ 1 Introduction
We have implemented the game Codenames  [^1] Source code (”codenames” directory): https://github.com/clembench/clembench as a challenging means (already investigated in psycholinguistic studies, see below) for evaluating certain language-use capabilities. In this game, a first player needs to provide to a second player a clue which singles out certain words within a larger set of words given to both players (see Figure 6 below for examples).
@@ 5.2 In-depth Analysis
Number of Targets, Guesses & Revealed: in Figure 2, we present the average number of words targeted and words guessed by selected models. We can see that high-performing models such as o3-mini and Deepseek-r1 generate at least 1-2 more words as targets and guesses in the beginning. Targeting and guessing more words in a single turn is the standard strategy in Codenames to win Spendlove and Ventura (2022), especially needed when playing against the mock opponent, which reveals one word at each turn. Models tend to guess fewer words than were targeted. For instance, o3-mini on average targets more than four words but guessed considerably fewer for the first turn, unlike Deepseek-r1, which targets and guesses an almost equal number of words. In Figure 7, we included the average number of target, guessed and revealed (where the guess is team word) words per model. We can see that only Deepseek-r1 exceeds the threshold of more than two words (2.2), while the rest have close values (1.5-1.9). It indicates that all models guess wrong words by revealing words from the opponent team or distractors, or even assassin words.
Success, Lose & Aborted Lose Rates:
@@
Typical Errors: To understand where models fail and how higher-performing models differ from lower ones, we analysed the most common errors, then categorised them and counted each occurrence, see Table 3.
@@
The differentiating factor in high-performing models is that hallucination and instruction following issues appear more rarely. For instance, the first error type, Target Hallucinated, refers to cases where Spymaster generates a clue and targets some words, but some of these do not exist on the board (as should be known to the model). In such cases, the GameMaster aborts that episode. Similarly, Guess Hallucinated is an error that occurs on the Field Operative side where it guesses a word that does not exist on the board. Mostly, Llama-3.1-405B and Llama-3.1-8B have another issue with guessing the correct number of words that the Spymaster indicates. They tend to guess more than the number of target words (note here: models can guess less but not more than target words). Lastly, the common issue, Guess is Clue, with low-performing models is that the guessed word is the same as the clue in many cases. It shows a lack of pragmatic reasoning for choosing unrevealed candidate words from available ones on a board. In all of these cases and some minor ones, e.g. tags such as “CLUE:”, “TARGETS:”, “GUESS:” are omitted, the GameMaster aborts the game because the rules are not followed. Such instruction-following issues happen mostly with Deepseek-r1 (see Table 5).